# FLIPHAT: Joint Differential Privacy for High Dimensional Sparse Linear Bandits

### v2

@@ 5.2 Regret Analysis
The first term in the upper bound $x_{\sf max}b_{\sf max}s^{*}\log d$ is intuitively the unavoidable part of regret, that is incurred during the initial part of the horizon when $\beta^{*}$ has not been estimated well enough. The second term is essentially two terms - $I_{\alpha}$ is the non-private statistical part of the regret while $J_{\alpha}(\epsilon,\delta)$ is the part arising due to the privacy constraint. The non-private term $I_{\alpha}$ is similar to the corresponding term in the regret bound for SLCB in (Chakraborty et al.,, 2023; Li et al.,, 2021). Furthermore, for $\alpha=0$, the upper bound in Theorem 5.7 matches with the problem-independent lower bound in Theorem 3.2 in terms of time horizon $T$ (up to logarithmic factors), context dimension $d$ and privacy parameters $\epsilon,\delta$. Now, we add the following remarks:
@@
1. Dependence on $d,s^{*},T$: The $({s^{*}}^{3}\log^{2}d)^{(1+\alpha)/2}$ term in the private part dominates the other parts (for fixed $\epsilon,\delta$). The dependence on $T$ is trickier: for the non-private part, for $\alpha=0$, the regret scales as $\sqrt{T}$, for $\alpha=1$ as $\log T$ and the effect fades as $\alpha\to\infty$. For the private part $J_{\alpha}$, as $\alpha\to 0$, $\Psi_{\alpha}(T)$ can be upper bounded by $\log^{3}T$, while the effect also fades here as $\alpha$ increases.
@@
2. Privacy trade-off for fixed $\alpha$: For any fixed level of $\alpha$, the regret bound explicates the threshold of the privacy parameter $\epsilon$, under which the cost of privacy dominates the non-private term. For example, with $\alpha=0$ (no margin condition), $J_{\alpha}$ dominates $I_{\alpha}$ when $\epsilon/\sqrt{\log(1/\delta)}<\log^{3}T\sqrt{(s^{*}\log d)/T}$. Similarly, for $\alpha=\infty$, the private part dominates when $\epsilon/\sqrt{\log(1/\delta)}<\sqrt{s^{*}\log d}$.
@@
3. Privacy vs. internal hardness: An interesting feature arising in the reget analysis is the mutual struggle between the hardness coming from the bandit instance’s arm-separation (through $\alpha$) and privacy constraint (through $\epsilon,\delta$). To understand this issue, let us denote $\bar{\epsilon}(\alpha)$ to be the threshold for $\epsilon$ under which the regret due to privacy dominates. We note that as $\alpha$ increases, $\bar{\epsilon}(\alpha)$ also increases, while keeping $\delta$ fixed. For example, $\bar{\epsilon}(\infty)=\sqrt{s^{*}\log d}\sqrt{\log(1/\delta)}$ and $\bar{\epsilon}(0)=\log^{3}T\sqrt{(s^{*}\log d)\log(1/\delta)/T}$. Thus, when the bandit instance is easy, i.e. for high $\alpha$, the hardness due to the privacy is dominant for a larger range of $\epsilon$ (analogously high $\bar{\epsilon}(\alpha)$). Conversely, when the bandit instance is itself hard, i.e. for low $\alpha$, the regret incurred due to this inherent difficulty quickly crosses the hardness due to privacy. As a result, even for smaller $\epsilon$, regret is dominated by the non-private part.
@@ 6 Numerical Experiments
Now, we present numerical experiments to illustrate the sub-linear dependence of regret on the ambient dimension $d$ under different privacy budgets $\epsilon$. In these experiments, we vary $d$ from 400 to about 2800. We set $s^{*}=10,K=3$ and $T=30,000$. We generate the contexts across time and arms from the Autoregressive Gaussian design in an i.i.d. fashion and with parameter $\phi=0.3$, i.e., $\Sigma_{ij}=0.3^{\left|i-j\right|}$ for $(i,j)\in[d]\times[d]$. Finally, the reward is generated according to the linear model (1) with Gaussian noise $N(0,0.1)$. All experiments were run on an online cluster with 16 GB RAM and 8 CPU cores.
Figure 1: Regret vs dimension for different privacy budgets.
This setup is the problem-independent setting, i.e. $\alpha=0$, as no prior information about the arm-separation is available. Figure 1 shows the final mean regret of FLIPHAT under $\delta=10^{-2}$ and $\epsilon\in\{0.8,2,5\}$ across 10 independent repetitions – a 95% confidence interval for the regret at each choice of $d$ is included in the figure as shaded bands around the mean regret lines. The dashed lines show that the regret depends almost logarithmically on the ambient dimension $d$. This matches with the theoretical results in Theorem 5.7, i.e., the cost of privacy in FLIPHAT scales as $O(\log d)$.
@@ 7 Conclusion
In this paper, we study the problem of high-dimensional sparse linear contextual bandit under joint differential privacy constraints. In particular, we establish novel minimax regret-bound results under JDP setting, demonstrating phase transition between the hardness of the problem depending on the privacy level. In addition, we propose a novel algorithm based on the doubling and forgetting trick that enjoys almost matching lower bounds in dimension, time horizon and privacy parameters under the JDP setting. To the best of our knowledge, this is the first JDP algorithm that achieves sharp rates in the regret bound under SLCB setting. It is worthwhile to mention that our algorithm is based on the N-IHT algorithm, for which we propose a refined convergence analysis under a more setting compared to the existing literature. As a consequence, our algorithm can be deployed in time sensitive high-risk tasks such as mobile-health applications for which high-dimensional information becomes a bottleneck from both computation and privacy point of view.
One natural extension of this work could be to explore the performance of FLIPHAT under the LDP setting, which is a stronger notion of privacy. Although, potentially the performance of FLIPHAT would be worse in the LDP setting compared to the JDP setting, it could be relevant in the field where the central server itself is not trusted. Another interesting direction could be to study the performance of FLIPHAT under an adversarial sparse LCB setting. It would also be an interesting future study to explore the lower bound using the margin condition to obtain bounds dependent on the hardness parameter $\alpha$ in the margin condition.
@@ Appendix A Background on differential privacy
  Initialize: $S=\emptyset$ and set $\xi=\lambda\frac{2\sqrt{3s\log(1/\delta)}}{\epsilon}$
  for $i=1,2,\dots,s$do
     Generate $\mathbf{w}_{i}=(w_{i1},\dots,w_{id})\overset{i.i.d.}{\sim}{\sf Lap}(\xi)$
     $j_{i}^{*}={\arg\max}_{j\in[d]\setminus S}|v_{j}|+w_{ij}$
     Update $S\leftarrow S\cup\{j_{i}^{*}\}$
  end for
  Set $\tilde{P}_{s}(v)=v_{S}$
  Generate $\tilde{\mathbf{w}}=(\tilde{w}_{1},\dots,\tilde{w}_{d})\overset{i.i.d.}{\sim}{\sf Lap}(\xi)$
  Return $\mathcal{P}_{s}(v;\epsilon,\delta,\lambda):=\tilde{P}_{s}(v)+\tilde{\mathbf{w}}_{S}$
@@ Appendix B Discussion on Assumptions
Assumption 5.3(a) about bounded contexts are standard in the bandit literature to obtain regret bounds independent of the scaling of the contexts. Assumption 5.3(b) says that the context distribution for each arm is a sub-Gaussian distribution with parameter $\vartheta$, this assumption does not require the distributions to have zero mean. This is indeed a very mild assumption on the context distribution and a broad class of distributions, such as truncated multivariate Gaussian and uniform on a compact set, enjoys such property. In comparison, several works in the literature on sparse contextual bandits (Oh et al.,, 2021; Kim and Paik,, 2019; Li et al.,, 2021) assume bounded $L_{2}$ norm, i.e. $\left\|X\right\|_{2}\leq L$, which in the high-dimensional setting is much stronger. Indeed $\left\|X\right\|_{2}\leq L$ implies $\left\|X\right\|_{\psi_{2}}\leq(L/\log 2)^{1/2}$ and in the example of truncated normal, $\left\|X\right\|_{2}=\Theta(\sqrt{d})$. Assumption 5.3(c) is an anti-concentration condition that ensures that the directions of the arms are well spread across every direction, thereby assisting in the estimation accuracy for $\beta^{*}$. Recent works (Li et al.,, 2021; Chakraborty et al.,, 2023) use similar assumptions to establish sharp estimation rates for $\beta^{*}$ under bandit settings. Assumption 5.3(d) imposes an upper bound on the maximum sparse eigenvalue of $\Sigma_{i}$ which is a common assumption in high-dimensional literature (Zhang and Huang,, 2008; Zhang,, 2010).
@@
Appendix C Regret Lower Bound - Details from Section 3
@@
Here, $P_{\mathbf{X}_{t},\mathbf{r}_{t}}$ is the joint distribution of the contexts $(x_{1}(t),\dots,x_{K}(t))\in\mathbb{R}^{Kd}$ and rewards $\mathbf{r}_{t}=(r_{1}(t),\dots,r_{K}(t))\in\mathbb{R}^{K}$ at time $t$, induced through $(x_{1}(t),\dots,x_{K}(t))\sim P_{\mathbf{X}}$ (with the marginal for $x_{i}(t)$ being $P_{i}$) and conditional on $x_{i}(t)$, $r_{i}(t)=x_{i}(t)^{\top}\beta^{*}+\epsilon_{i}(t)$. $\epsilon_{i}(t)\sim P_{\epsilon}$ independently for all $i,t$. A particular choice of $\beta^{*},P_{\mathbf{X}}$ and $P_{\epsilon}$ lead to a particular bandit instance $P_{\mathbf{X},\mathbf{r}}\in\mathcal{B}_{\beta^{*}}$. To control for the scaling in the contexts and the conditional rewards, we assume that for $(x_{1},\dots,x_{K})\sim P_{\mathbf{X}}$, the marginal covariance of each $x_{k}$ has eigenvalues bounded by $\sigma_{X,\min}^{2}$ and $\sigma_{X,\max}^{2}$. We also assume that $P_{\epsilon}$ is a sub-Gaussian distribution with variance-proxy $\sigma^{2}$. The dependence on these parameters will be denoted by $\mathcal{B}_{\beta^{*},\sigma_{X},\sigma}$ in the space of joint distributions of contexts and rewards. However, given a particular bandit instance, the learner does not get to observe all this data. This is where the policy comes in.
@@ D.1 Results on Peeling
$\overset{\eta\underline{\kappa}<1}{\leq}\left(1-\eta\underline{\kappa}+\frac{\eta}{c_{3}}\right)\left\|\theta^{m}_{I_{m}}-\theta^{*}\right\|_{2}^{2}+\left(\eta^{2}+\eta c_{3}(1-\eta\underline{\kappa})^{2}\right)\left\|\nabla\mathcal{L}(\theta^{*})_{I_{m}}\right\|_{2}^{2}$
@@
We note that by algorithm, $\theta^{m}$ is utmost $s$ sparse and we would chosen $s\asymp\kappa^{2}s^{*}$ (made precise later). Inequality ()(**)( ∗ ∗ ) above now follows by using the RSC parameter $\underline{\kappa}$ from (9) and arguing as
@@
Now, with the choices $\eta=1/2\bar{\kappa},c_{1}=56\kappa-3,c_{2}=1/8(14\kappa-1),c=1$ and $s=16(14\kappa-1)(112\kappa-7)s^{*}$, we have for suitably large constants $C_{1},C_{2}>0$
@@
$b_{3}$ $\leq C_{2}\frac{\kappa}{\underline{\kappa}^{2}}s^{*}.$
@@
$\left\|\theta^{m+1}-\theta^{*}\right\|_{2}^{2}\leq\rho\left\|\theta^{m}-\theta^{*}\right\|_{2}^{2}+C_{1}\kappa\underbrace{\left(\left\|\tilde{\mathbf{w}}_{S^{m}}^{(m)}\right\|_{2}^{2}+\sum_{i\in[s]}\left\|\mathbf{w}_{i}^{(m)}\right\|_{\infty}^{2}\right)}_{\mathbf{W}_{m}}+C_{2}\frac{\kappa}{\underline{\kappa}^{2}}s^{*}\left\|\nabla\mathcal{L}(\theta^{*})\right\|_{\infty}^{2}.$
@@
$\left\|\theta^{M}-\theta^{*}\right\|_{2}^{2}$ $\leq\rho^{M}\left\|\theta^{0}-\theta^{*}\right\|_{2}^{2}+C_{1}\kappa\sum_{m=0}^{M-1}\rho^{m}\mathbf{W}_{m}+C_{2}\frac{\kappa}{\underline{\kappa}^{2}}s^{*}\left\|\nabla\mathcal{L}(\theta^{*})\right\|_{\infty}^{2}\sum_{m=0}^{M-1}\rho^{m}$
$\leq 2\rho^{M}b_{\sf max}^{2}+C_{1}\kappa\left(\max_{m}\mathbf{W}_{m}\right)\frac{1-\rho^{M}}{1-\rho}+C_{2}\frac{\kappa}{\underline{\kappa}^{2}}s^{*}\left\|\nabla\mathcal{L}(\theta^{*})\right\|_{\infty}^{2}\frac{1-\rho^{M}}{1-\rho}$ (13)
@@
$\left\|\theta^{M}-\theta^{*}\right\|_{2}^{2}\leq\frac{1}{n}+C\kappa^{2}\bar{W}+C^{\prime}\frac{\kappa^{2}}{\underline{\kappa}^{2}}s^{*}\bar{g}$ (14)

### v3

@@ 5.2 Regret Analysis
The first term in the upper bound $x_{\sf max}b_{\sf max}s^{*}\log d$ is intuitively the unavoidable part of regret, that is incurred during the initial part of the horizon when $\beta^{*}$ has not been estimated well enough. The second term is essentially two terms - $I_{\alpha}$ is the non-private statistical part of the regret while $J_{\alpha}(\epsilon,\delta)$ is the part arising due to the privacy constraint. The non-private term $I_{\alpha}$ is similar to the corresponding term in the regret bound for SLCB in (Chakraborty et al.,, 2023; Li et al.,, 2021). Furthermore, for $\alpha=0$, the upper bound in Theorem 5.6 matches with the problem-independent lower bound in Theorem 3.2 in terms of time horizon $T$ (up to logarithmic factors), context dimension $d$ and privacy parameters $\epsilon,\delta$. Below, we state the implications.
@@
1. Dependence on $d,s,s^{*},T$: The term in $J(\epsilon,\delta)$ scaling as $\{(s+s^{*})^{3}\log^{2}d\}^{(1+\alpha)/2}$ dominates the non-private term $I_{\alpha}$ (for fixed $\epsilon,\delta$). The dependence on $T$ is trickier: for the non-private part, for $\alpha=0$, the regret scales as $\sqrt{T}$, for $\alpha=1$ as $\log T$ and the effect fades as $\alpha\to\infty$. For the private part $J_{\alpha}$, as $\alpha\to 0$, $\Psi_{\alpha}(T)$ can be upper bounded by $\log^{3}T$, while the effect fades as $\alpha$ increases.
@@
2. Privacy trade-off for fixed $\alpha$: For any fixed level of $\alpha$, the regret bound explicates the threshold of the privacy parameter $\epsilon$, under which the cost of privacy dominates the non-private term. For example, with $\alpha=0$ (no margin condition), $J_{\alpha}$ dominates $I_{\alpha}$ when $\epsilon/\sqrt{\log(1/\delta)}<\log^{3}T\sqrt{\{(s+s^{*})\log d\}/T}$. Similarly, for $\alpha=\infty$, the private part dominates when $\epsilon/\sqrt{\log(1/\delta)}<\sqrt{(s+s^{*})\log d}$.
@@
3. Privacy vs. internal hardness: An interesting feature arising in the reget analysis is the mutual struggle between the hardness coming from the bandit instance’s arm-separation (through $\alpha$) and privacy constraint (through $\epsilon,\delta$). To understand this issue, let us denote $\bar{\epsilon}(\alpha)$ to be the threshold for $\epsilon$ under which the regret due to privacy dominates. We note that as $\alpha$ increases, $\bar{\epsilon}(\alpha)$ also increases, while keeping $\delta$ fixed. For example, $\bar{\epsilon}(\infty)=\sqrt{(s+s^{*})\log d}\sqrt{\log(1/\delta)}$ and $\bar{\epsilon}(0)=\log^{3}T\sqrt{\{(s+s^{*})\log d\}\log(1/\delta)/T}$. Thus, when the bandit instance is easy, (high $\alpha$), the hardness due to the privacy is dominant for a larger range of $\epsilon$ (analogously high $\bar{\epsilon}(\alpha)$). Conversely, when the bandit instance is itself hard (low $\alpha$), the regret incurred due to this inherent difficulty quickly crosses the hardness due to privacy. As a result, even for smaller $\epsilon$, regret is dominated by the non-private part.
4. Achieving minimax optimal bound: The regret bound in Theorem 5.6 crucially depends on the quantity $(s+s^{*})$. Therefore, the choice of $s$ is very important to avoid the curse of dimensionality. In a high-dimensional sparse setting, $s^{*}$ is typically very small, i.e., $s^{*}\ll d$. Therefore, setting $s$ in the order of $o(d)$ is also a valid choice, and an improved dependence on $d$ can be achieved. Moreover, under a constant order knowledge of $s^{*}$, the tuning parameter $s$ can be chosen as $s\asymp\kappa^{2}s^{*}$ which leads to a near-optimal dependence in $s^{*}$. In addition, numerical experiments in Appendix F.3 show that the regret only scales polynomially in $s>s^{*}$ as shown in Theorem 5.6.
@@ 6 NUMERICAL EXPERIMENTS
We present an empirical study of the performance of FLIPHAT through several numerical experiments. All experiments were run locally on Macbook Pro with 16 GB RAM and 1.4 GHz Quad-Core Intel Core i5. Both of the experiments are setup in the problem-independent setting $\alpha=0$, and the plots show both the mean and $95\%$-CI for the mean (as colored bands) over 60 repetitions of each experiment. For space constraint, we present two of these experiments here, and postpone the additional details and studies to the Appendix F.
In the first experiment, we test the time evolution of regret of FLIPHAT, and illustrate the regret against time $T$ in Figure 1(Left) in the case where $d=400,s^{*}=5,K=3$ up to $T=20000$ for different choices of privacy parameter $\epsilon\in\{0.5,1,2,5,10\},\delta=0.01$. We also include random bandit (taking random action at each time), which corresponds to $\epsilon=0$ (full privacy, no effective utility), and also the sparsity agnostic bandit (Oh et al.,, 2021), which corresponds to $\epsilon\to\infty$, i.e. without any privacy guarantee, as the baselines. The contexts are drawn i.i.d. from $\mathcal{N}(0,\Sigma)$, where $\Sigma_{ij}=0.1^{|i-j|}$ (autoregressive design), with observation noise being a centered Gaussian with $\sigma=0.1$. For the second experiment, we illustrate the effect of context dimension $d$ on the regret suffered by FLIPHAT at $T=10000$ for 4 different choices of $\epsilon$ (Figure 1-Right) in a setting similar to the first experiment, while $d$ takes 12 equi-spaced values from 400 to 4000.
Results. Figure 1 (Left) demonstrates the sublinear behavior of FLIPHAT’s regret that approaches the sparsity agnostic bandit’s regret for larger $\epsilon$. Figure 1 (Right) shows that $R(T)$ with $T=10000$ is nearly linear in $\log d$, where the slope $\beta$ (in the legend) is computed by fitting a linear model. This matches with the theoretical results in Theorem 5.6, i.e., the cost of privacy in FLIPHAT scales as $O(\log d)$.
Figure 1: (Left) Regret vs $T$ for different privacy level $\epsilon$, (Right) Regret at $T=10000$ vs $\log(d)$ for different $\epsilon$
@@ 7 CONCLUSION
In this paper, we study the SLCBs under JDP constraints. We establish two minimax regret lower bounds under JDP settings, demonstrating phase transition between the hardness of the problem depending on the privacy level. We propose FLIPHAT, which is the first algorithm to achieves sharp regret-privacy trade-off in SLCBs. FLIPHAT is based on the N-IHT algorithm, for which we propose a refined convergence analysis under a general setting than the existing literature.
An interesting direction would be to study FLIPHAT under an adversarial sparse LCB setting. It would also be an interesting future study to explore the lower bound using the margin condition and obtain bounds dependent on the parameter $\alpha$ in the margin condition.
@@ Appendix A BACKGROUND ON DIFFERENTIAL PRIVACY
1:  Initialize: $S=\emptyset$ and set $\xi=\lambda\frac{2\sqrt{3s\log(1/\delta)}}{\epsilon}$
2:  for $i=1,2,\dots,s$do
3:     Generate $\mathbf{w}_{i}=(w_{i1},\dots,w_{id})\overset{i.i.d.}{\sim}{\sf Lap}(\xi)$
4:     $j_{i}^{*}={\arg\max}_{j\in[d]\setminus S}|v_{j}|+w_{ij}$
5:     Update $S\leftarrow S\cup\{j_{i}^{*}\}$
6:  end for
7:  Set $\tilde{P}_{s}(v)=v_{S}$
8:  Generate $\tilde{\mathbf{w}}=(\tilde{w}_{1},\dots,\tilde{w}_{d})\overset{i.i.d.}{\sim}{\sf Lap}(\xi)$
9:  Return $\mathcal{P}_{s}(v;\epsilon,\delta,\lambda):=\tilde{P}_{s}(v)+\tilde{\mathbf{w}}_{S}$
@@ Appendix B DISCUSSION ON ASSUMPTIONS
Assumption 5.2(a) about bounded contexts are standard in the bandit literature to obtain regret bounds independent of the scaling of the contexts. Assumption 5.2(b) says that the context distribution for each arm is a sub-Gaussian distribution with parameter $\vartheta$, this assumption does not require the distributions to have zero mean. This is indeed a very mild assumption on the context distribution and a broad class of distributions, such as truncated multivariate Gaussian and uniform on a compact set, enjoys such property. In comparison, several works in the literature on sparse contextual bandits (Oh et al.,, 2021; Kim and Paik,, 2019; Li et al.,, 2021) assume bounded $L_{2}$ norm, i.e. $\left\|X\right\|_{2}\leq L$, which in the high-dimensional setting is much stronger. Indeed $\left\|X\right\|_{2}\leq L$ implies $\left\|X\right\|_{\psi_{2}}\leq(L/\log 2)^{1/2}$ and in the example of truncated normal, $\left\|X\right\|_{2}=\Theta(\sqrt{d})$. Assumption 5.2(c) is an anti-concentration condition that ensures that the directions of the arms are well spread across every direction, thereby assisting in the estimation accuracy for $\beta^{*}$. Recent works (Li et al.,, 2021; Chakraborty et al.,, 2023) use similar assumptions to establish sharp estimation rates for $\beta^{*}$ under bandit settings. Assumption 5.2(d) imposes an upper bound on the maximum sparse eigenvalue of $\Sigma_{i}$ which is a common assumption in high-dimensional literature (Zhang and Huang,, 2008; Zhang,, 2010). Below we define the minimum and maximum sparse eigenvalue conditions formally.
Definition B.1 (Sparse Riesz Condition).
$\phi_{\min}(s;A):=\inf_{u:u\neq 0,\left\|u\right\|_{0}\leq s}u^{\top}Au/\left\|u\right\|_{2}^{2},\quad\phi_{\max}(s;A):=\sup_{u:u\neq 0,\left\|u\right\|_{0}\leq s}u^{\top}Au/\left\|u\right\|_{2}^{2}\,.$ (4)
@@
Appendix C REGRET LOWER BOUND: DETAILS FROM SECTION 3
@@
Here, $P_{\mathbf{X}_{t},\mathbf{r}_{t}}$ is the joint distribution of the contexts $(x_{1}(t),\dots,x_{K}(t))\in\mathbb{R}^{Kd}$ and rewards $\mathbf{r}_{t}=(r_{1}(t),\dots,r_{K}(t))\in\mathbb{R}^{K}$ at time $t$, induced through $(x_{1}(t),\dots,x_{K}(t))\sim P_{\mathbf{X}}$ (with the marginal for $x_{i}(t)$ being $P_{i}$) and conditional on $x_{i}(t)$, $r_{i}(t)=x_{i}(t)^{\top}\beta^{*}+\epsilon_{i}(t)$. $\epsilon_{i}(t)\sim P_{\epsilon}$ independently for all $i,t$. A particular choice of $\beta^{*},P_{\mathbf{X}}$ and $P_{\epsilon}$ lead to a particular bandit instance $P_{\mathbf{X},\mathbf{r}}\in\mathcal{B}_{\beta^{*}}$. To control for the scaling in the contexts and the conditional rewards, we assume that for $(x_{1},\dots,x_{K})\sim P_{\mathbf{X}}$, the marginal covariance of each $x_{k}$ has eigenvalues bounded by $\sigma_{X,\min}^{2}$ and $\sigma_{X,\max}^{2}$. We also assume that $P_{\epsilon}$ is a sub-Gaussian distribution with variance-proxy $\sigma^{2}$. The dependence on these parameters will be denoted by $\mathcal{B}_{\beta^{*},\sigma_{X},\sigma}$ in the space of joint distributions of contexts and rewards.. However, given a particular bandit instance, the learner does not get to observe all this data. This is where the policy comes in.
@@ D.1 Results on Peeling
$\overset{\eta\underline{\kappa}<1}{\leq}\left(1-\eta\underline{\kappa}+\frac{\eta}{c_{3}}\right)\left\|\theta^{m}_{I_{m}}-\theta^{*}\right\|_{2}^{2}+\left(\eta^{2}+\eta c_{3}(1-\eta\underline{\kappa})^{2}\right)\left\|\nabla\mathcal{L}(\theta^{*})_{I_{m}}\right\|_{2}^{2}$ (**)
@@
We note that by algorithm, $\theta^{m}$ is at most $s$ sparse and we would choose $s\gtrsim\kappa^{2}s^{*}$ (made precise later). Inequality ()(**)( ∗ ∗ ) above now follows by using the RSC parameter $\underline{\kappa}$ from (10) and arguing as before.
@@
Now, with the choices $\eta=1/2\bar{\kappa},c_{1}=56\kappa-3,c_{2}=1/8(14\kappa-1),c=1$ and $s\geq 16(14\kappa-1)(112\kappa-7)s^{*}$, we have for suitably large constants $C_{1},C_{2}>0$
@@
$b_{3}$ $\leq C_{2}\frac{\kappa}{\underline{\kappa}^{2}}(s+s^{*}).$
@@
$\left\|\theta^{m+1}-\theta^{*}\right\|_{2}^{2}\leq\rho\left\|\theta^{m}-\theta^{*}\right\|_{2}^{2}+C_{1}\kappa\underbrace{\left(\left\|\tilde{\mathbf{w}}_{S^{m}}^{(m)}\right\|_{2}^{2}+\sum_{i\in[s]}\left\|\mathbf{w}_{i}^{(m)}\right\|_{\infty}^{2}\right)}_{\mathbf{W}_{m}}+C_{2}\frac{\kappa}{\underline{\kappa}^{2}}(s+s^{*})\left\|\nabla\mathcal{L}(\theta^{*})\right\|_{\infty}^{2}.$
@@
$\left\|\theta^{M}-\theta^{*}\right\|_{2}^{2}$ $\leq\rho^{M}\left\|\theta^{0}-\theta^{*}\right\|_{2}^{2}+C_{1}\kappa\sum_{m=0}^{M-1}\rho^{m}\mathbf{W}_{m}+C_{2}\frac{\kappa}{\underline{\kappa}^{2}}(s+s^{*})\left\|\nabla\mathcal{L}(\theta^{*})\right\|_{\infty}^{2}\sum_{m=0}^{M-1}\rho^{m}$
$\leq 2\rho^{M}b_{\sf max}^{2}+C_{1}\kappa\left(\max_{m}\mathbf{W}_{m}\right)\frac{1-\rho^{M}}{1-\rho}+C_{2}\frac{\kappa}{\underline{\kappa}^{2}}(s+s^{*})\left\|\nabla\mathcal{L}(\theta^{*})\right\|_{\infty}^{2}\frac{1-\rho^{M}}{1-\rho}$ (14)
@@
$\left\|\theta^{M}-\theta^{*}\right\|_{2}^{2}\leq\frac{1}{n}+C\kappa^{2}\bar{W}+C^{\prime}\frac{\kappa^{2}}{\underline{\kappa}^{2}}(s+s^{*})\bar{g}$ (15)